OpenStudy (anonymous):

degree measure of an angle is 40° less than 3 times the degree measure of its supplementary angle, what is the degree measure of the angle? A. 50° B. 55° C. 125° D. 130°

OpenStudy (mathstudent55):

Do you know what supplementary angles are?

OpenStudy (anonymous):

180 degrees

OpenStudy (kash_thesmartguy):

Angles which have a sum of 180degrees!

OpenStudy (mathstudent55):

Correct. Two angles whose measures add up to 180 degrees.

OpenStudy (mathstudent55):

Let's call for now the angles x and y. From the fact they are supplementary, we know: x + y = 180

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

90 plus 90

OpenStudy (anonymous):

are u there

OpenStudy (mathstudent55):

90 plus 90 is an example of two supplementary angles Now let's see what our supplementary angles measure.

OpenStudy (anonymous):

so the answer would be c then

OpenStudy (mathstudent55):

Once again, the angles measure x and y. x + y = 180 Let's solve this equation for y: y = 180 -x That means the two angles measure x and 180 - x

OpenStudy (mathstudent55):

Now we use the given info. "an angle is 40° less than 3 times the degree measure of its supplementary angle" x = 3(180 - x) - 40 3(180 - x) - 40 + 180 - x = 180 540 - 3x - 40 - x = 0 -4x + 500 = 0 -4x = -500 x = 125

OpenStudy (mathstudent55):

Correct. The answer is C. 125 deg.